OpenStudy (anonymous):

can some one help me A truck can be rented from company A for $110 a day plus $0.40 per mile. Company b charges $70 a day plus $0.80 per mile to rent the same truck. How many miles must be driven in a day to make the rental cost company a better deal than company B.?

OpenStudy (kainui):

So lets break it down a little. If it's $110 a day, then how much would it be for 2 days? Well since the first day was 110 and the second day was 110 then it must be 220. If you go a third day it's 330. Now you're starting to see the pattern here? The number of days multiplied by the price is the total price. $110*d=part of the total price Now see if you can make up another piece to add to that based on the total miles.

OpenStudy (kainui):

When it comes down to it this is an easy question for many people and they have no problems giving you the answer, but if you don't try and work with me while I'm here you'll just fail your tests and not learn anything.

OpenStudy (eyust707):

... so from what @Kainui said we can begin to write an equation that represents the total price of company A. So far we know that each day costs 110. 110 * the number of days = how much we get charged per day Now we need to add in a term that represents how much we will be charged per mile.

OpenStudy (eyust707):

We can do the same thing per mile.. How much is it per mile for company A?

OpenStudy (anonymous):

sorry i had to do something this question is not from a test or home work

OpenStudy (eyust707):

sweet so your doing it for fun just to learn it?

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

awesome! I will help you learn how to do this

OpenStudy (anonymous):

thank you

OpenStudy (eyust707):

Lets reread the first part of the question. "A truck can be rented from company A for $110 a day plus $0.40 per mile. "

OpenStudy (eyust707):

The goal is to write an equation for this sentence

OpenStudy (anonymous):

okay

OpenStudy (eyust707):

"110 per day plus .40 per mile"

OpenStudy (eyust707):

that is the same as: 100 * days + .40 * miles

OpenStudy (anonymous):

okay

OpenStudy (eyust707):

sry 110 *days + .40 * miles

OpenStudy (eyust707):

simplified lets say: Cost of A = 110 * d + .40 * m

OpenStudy (eyust707):

do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

great! now it is your turn.

OpenStudy (eyust707):

In the same manner write the following as a equation: Company b charges $70 a day plus $0.80 per mile

OpenStudy (anonymous):

B=70*d+70*

OpenStudy (eyust707):

close

OpenStudy (eyust707):

B=70*d + .8 * m

OpenStudy (anonymous):

oh okay

OpenStudy (eyust707):

Now looking at both: A = 110*d + .40*m B = 70*d + .8 * m

OpenStudy (anonymous):

okay

OpenStudy (eyust707):

So just looking at it from a logical stand point real quick, A costs more per day but is cheaper per mile and B costs less per day but is more expensive per mile.

OpenStudy (anonymous):

yes

OpenStudy (eyust707):

If you were going to rent a truck and drive a really long distance which company would you use?

OpenStudy (anonymous):

B

OpenStudy (eyust707):

but B charges a lot per mile!

OpenStudy (eyust707):

for A you have to pay a lot at first... but the miles are really cheap

OpenStudy (anonymous):

oh sorry i just saw the question i meant to say a

OpenStudy (eyust707):

sweet! so you see where this question is headed

OpenStudy (eyust707):

how many miles do I have to drive in order to make it worth it to go with A and not B?

OpenStudy (eyust707):

Well first of all the question says it is for 1 day. So we can set d = 1

OpenStudy (eyust707):

Now we have: A = 110 + .40*m B = 70 + .8 * m

OpenStudy (anonymous):

okay

OpenStudy (eyust707):

Would you agree with me that there exists a certain number of miles where if I went with A or B it wouldnt matter because they would cost the same?

OpenStudy (anonymous):

yes i agree

OpenStudy (eyust707):

Great! So if we drive more than that amount A is the better choice and less then B is the better choice.

OpenStudy (eyust707):

Lets find that amount!

OpenStudy (eyust707):

To do that we need to solve for the number of miles that makes the total cost the same

OpenStudy (anonymous):

okay

OpenStudy (eyust707):

A = 110 + .40*m B = 70 + .8 * m We want A and B to be equal then we want to solve for m

OpenStudy (eyust707):

or: 110 + .40*m = 70+ .8 * m

OpenStudy (eyust707):

each of those equations represents the cost and when the cost is equal we have the cool point where it doesnt matter which one we go to

OpenStudy (eyust707):

can you solve for m? 110 + .40*m = 70+ .8 * m

OpenStudy (anonymous):

okay i have 181.2

OpenStudy (eyust707):